OpenStudy (anonymous):

please can you show me how to simplify the following expression sqrt sign -7 then sqrt sign -105

OpenStudy (anonymous):

whenever you're taking the square root of some negative number, write it out "i" on the outside of the radical and change the radicand to positive.

OpenStudy (anonymous):

can you please simplify and show me the answer in detail?

OpenStudy (anonymous):

i then sqrt 7 then i sqrt 105 would be the answer?

OpenStudy (anonymous):

ok.. for example... sqrt(-7): \[\large \sqrt {-7} = i \times \sqrt {7} = \sqrt {7}i \]

OpenStudy (anonymous):

then the other would be the same other then it is 105 is that correct?

OpenStudy (anonymous):

yes..:)

OpenStudy (anonymous):

ok then i do understand it seeings how you made it. thank you, appreciate it:)

OpenStudy (anonymous):

glad to have helped... & yw... :)

OpenStudy (anonymous):

√ -7 = i * √7 = √7i and √-105 = i * √105 = √105i would the "and" be a * sign in the middle?